OpenStudy (anonymous):

What is the third–degree polynomial function such that f(0) = –24 and whose zeros are 1, 2, and 3?

OpenStudy (anonymous):

f(0)=4(x-1)(x-2)(x-3)=-24

OpenStudy (anonymous):

you can convert that into a cube function

OpenStudy (anonymous):

General \[f(x)=a(x-1)(x-2)(x-3) \\f(0)=a(0-1)(0-2)(0-3)=-24 \\ \therefore \ a = 4\]

OpenStudy (anonymous):

f(x) = 4x3 – 24x2 + 44x – 24 f(x) = 2x3 – 24x2 + 22x – 24 f(x) = 4x3 + 24x2 – 44x – 24 f(x) = 2x3 + 24x2 – 22x – 24 -these are the options .

OpenStudy (anonymous):

a